Question 7 Table 8-2 Vanadia Company acquired a $40,000 machine on January 1, 20X9. The machine is estimated to have a useful life of 5 years, and a residual value of $4,000. For unit depreciation purposes, the machine is expected to produce 500,000 units. Referring to Table 8-2, if Vanadia Company uses unit depreciation, and the company produces 80,000 units in 20X9, what will be the depreciation expense for 20X9? $8,000 $6,400 $8,800 $5,760 $7,200
Answer #1

Solution:

Depreciation expense per unit = (40000 - 4000) / 500000 = $0.072

Depreciation expense for 20x9 = 0.072*80000 = $5,760